Cd=200+0.5Y-500r
Id=200-500r
L=0.5Y-250(r+à ƒ ƒ  €e)
à ƒ ƒ  €e=0
G=150
M=4900
Y=100

Derive the equilibrium levels of the real interest rates r, desired consumption Cp,desired investment Io, and the price level p ,L is money demand equation and M is money demand equation
1
Expert's answer
2021-10-20T10:16:59-0400

You

21:38

Y = C + I + G

Y = 200 + 0.25YD + 150 + 0.25Y – 1000i + 250

Y = 200 + 0.25(Y – T) +150 + 0.25Y – 1000i + 250

Y = 200 + 0.25Y – 0.25T + 150 + 0.25Y – 1000i + 250

Y = 200 + 0.25Y – 0.25(200) + 150 + 0.25Y – 1000i + 250

Y = 200 + 0.25Y – 50 + 150 + 0.25Y – 1000i + 250

Y = 0.5Y + 550 – 1000i

0.5Y = 550 – 1000i

Y = 1100 – 2000i : Which is the IS relation.